Find the probability of the following: (Be sure to show all work)

I have a part time job in food industry and probability is very important on selling food items everyday. We sold 24 hamburger baskets and 16 had cheese on them. 6 of the hamburger baskets were ordered to go. 10 of the cheeseburger baskets were ordered to go. If a hamburger basket or cheeseburger basket were selected at random, find the probability of A) It was a Hamburger basket to go B) It was a cheeseburger basket that was ordered to go.
Answer #1

probability of a Hamburger basket to go = (total hamburger order)/ total baskets = 6/(24+16) =  6/40

probability of a cheeseburger basket to go = (total cheeseburger order)/ total baskets = 10/(24+16) = 10/40